OpenStudy (anonymous):

Systems of Equations, please help =) Solve without matrix 5x-y+z=4; x+2y-2=5; 2x+3y-3=5 I have the answer, but have tried multiple times and cannot get to it.

OpenStudy (anonymous):

WAIT that's the wrong equation!! It's 5x-y+Z=4, x+2y-Z=5, 2x+3y+3Z=5

OpenStudy (campbell_st):

so which letter do you want to eliminate... y or z..?

OpenStudy (anonymous):

Whichever is easier.

OpenStudy (campbell_st):

ok... so we'll eliminate z you'll end up with 2 equations in 2 unknowns so add equations 1 and 2 5x -y + z = 4 x + 2y - z = 5 -------------- 6x + y = 9 equation 4 now multiply equation 2 by 3 and add equation 3 3x + 6y -3z = 15 2x + 3y + 3z = 5 ----------------- 5x + 9y = 20 equation 5 easiest thing to do now is to multiply equation 4 by 9 and subtract equation 5 which will eliminate y 54x + 9y = 81 - 5x - 9y = -20 -------------- 49x = 61

OpenStudy (campbell_st):

you'll need to check the calculations

OpenStudy (anonymous):

So x=(61/49)?

OpenStudy (campbell_st):

the unlucky thing about the question you posted is that the answers are all ugly fractions... but I hope it helps you should be able to get x then substitute it into either equation 4 or 5 to find y then substitute the values of x and y into any of the original equations to find x hope it makes sense

OpenStudy (campbell_st):

that's x... very ugly

OpenStudy (anonymous):

Huh. The answer is x=1, y=3, z=2...

OpenStudy (campbell_st):

well you need to check the equations 1st seems ok... 5 - 3 + 2 = 4 2nd 1 + 6 - 2 = 5 3rd 2 + 9 + 6 = 17 you have 5 can you check the 3rd equation in the repost....

OpenStudy (campbell_st):

so it looks like the last equation should be 2x + 3y - 3z = 5

OpenStudy (anonymous):

Yes. Oops!! So sorry

OpenStudy (campbell_st):

ok... so equation 4 is fine... now multiply equation 1 by 3 and then add equation 3\\ 15x - 3y + 3z = 12 2x + 3y + 3z = 5 ------------------ 17x = 17 so now that makes sense... find x, substitute it into equation 4 to find y then substitute the values into any of the original equations to find z hope it now makes sense.

OpenStudy (campbell_st):

oops should read -3z = 5

OpenStudy (anonymous):

yes!! thank you so much!!